# Self-Supervised Frameworks for Speaker Verification via Bootstrapped Positive Sampling

### v2

@@ Abstract
Recent developments in Self-Supervised Learning (SSL) have demonstrated significant potential for Speaker Verification (SV), but closing the performance gap with supervised systems remains an ongoing challenge. Standard SSL frameworks rely on anchor-positive pairs extracted from the same audio utterances. Hence, positives have channel characteristics similar to those of their corresponding anchors, even with extensive data-augmentation. Therefore, this positive sampling strategy is a fundamental limitation as it encodes too much information regarding the recording source in the learned representations. This article introduces Self-Supervised Positive Sampling (SSPS), a bootstrapped technique for sampling appropriate and diverse positives in SSL frameworks for SV. SSPS samples positives close to their anchor in the representation space, under the assumption that these pseudo-positives belong to the same speaker identity but correspond to different recording conditions. This method demonstrates consistent improvements in SV performance on VoxCeleb benchmarks when implemented in major SSL frameworks, such as SimCLR, SwAV, VICReg, and DINO. Using SSPS, SimCLR, and DINO achieve 2.57% and 2.53% EER on VoxCeleb1-O. SimCLR yields a 58% relative reduction in EER, getting comparable performance to DINO with a simpler training framework. Furthermore, SSPS lowers intra-class variance and reduces channel information in speaker representations while exhibiting greater robustness without data-augmentation.
@@ I Introduction
Speaker Recognition (SR) identifies the speaker’s identity in an audio speech utterance. The main task in the SR field is Speaker Verification (SV), which aims to determine whether two speech utterances were spoken by the same speaker. To achieve this task, SR systems aim to define representations that maximize inter-speaker distances and minimize intra-speaker variances while being robust to extrinsic variabilities (e.g., noise, environment, recording, and channel conditions).
@@
With the emergence of deep learning, traditional machine learning methods such as i-vectors [dehak2011IVector] have been outperformed by DNN models such as the x-vectors [snyder2018XVectors], architectures based on ResNet [chung2020DefenceMetricLearningSR], and more recently, the ECAPA-TDNN [desplanques2020ECAPATDNN] model. These methods are trained to match input speech utterances to their respective speaker identities in a supervised manner on a large corpus of annotated speech segments [chung2018VoxCeleb2]. Indeed, the performance of deep learning methods scales as the amount of training data increases. However, this training paradigm represents a notable constraint due to the scarcity and expense of getting annotated training samples.
@@
Self-Supervised Learning (SSL) methods have emerged as a solution to this bottleneck by learning meaningful representations from the input data. SSL has been shown to enhance the scalability potential of models by leveraging the abundance of unlabeled data available. Various SSL frameworks have been proposed in Computer Vision (CV) following the joint-embedding architecture, which considers an anchor and a positive sample derived from different augmentations of the same input, thereby representing the same high-level information. Contrastive learning, adopted by SimCLR [chen2020SimCLR] and MoCo [he2020MoCo], aims to maximize the similarity of positive pairs while minimizing the similarity of negative pairs sampled from the current batch or a memory queue. The effect of class collision, occurring when randomly sampled negatives belong to the same class as the anchor-positive pair, is considered negligible on the training convergence. Clustering-based methods, such as DeepCluster [caron2018DeepCluster] and SwAV [caron2020SwAV], learn representations by grouping similar representations into clusters and using the resulting assignments as a supervisory signal. Information maximization techniques like Barlow Twins [zbontar2021BarlowTwins] and VICReg [bardes2022VICReg] apply constraints on the learned embeddings space to enforce invariance between the anchor and the positive embeddings while explicitly avoiding collapse (i.e., non-informative representations). Finally, methods based on self-distillation, such as BYOL [grill2020BYOL] and DINO [caron2021DINO], leverage knowledge distillation and its student-teacher paradigm, where a student network is trained to match the teacher network embeddings.
@@
These approaches have also been successfully extended to the downstream task of SV. Most methods are based on contrastive learning [huh2020APAAT, zhang2021SimCLR, xia2021MoCo, lepage2022LabelEfficientSSLSV, lepage2024AdditiveMarginSSLSV] and knowledge distillation, as the DINO framework is currently state-of-the-art for many downstream tasks, including SV [zhang2022C3DINO, chen2023RDINO, han2024CADINO, cho2022DINO, heo2022DINOCurriculum]. In the context of SV, SSL relies on the principle that the anchor and the positive belong to the same speaker, as both segments are extracted from the same speech utterance (Figure 1a). Moreover, extensive data-augmentation is applied to both these segments to learn robust representations against extrinsic variabilities (environmental noise, reverberation, mismatching recording devices, etc.).
@@
Nonetheless, it has been shown that data-augmentation is insufficient to overcome the effect of SSL same-utterance positive sampling, which encodes various channel characteristics in speaker representations, yielding high variance within same-speaker representations. Several techniques have been explored to address this issue in SV: AP+AAT [huh2020APAAT] introduced an adversarial loss to penalize the model from learning channel information; i-mix [kang2022LMix] is a data-driven augmentation strategy that interpolates training utterances to make the model focus on the main distinguishing factor between them; DPP [tao2023DPP] finds diverse positive pairs using another modality by cross-referencing speech and face data; CA-DINO [han2024CADINO] clusters the embeddings at the end of training to sample positives from the same class as the anchor. Finally, alternative positive sampling strategies have been proposed in CV, such as NNCLR [dwibedi2021NNCLR] and GPS-SSL [feizi2024GPSSSL], which find positives in the latent space using the nearest neighbors search.
@@
A new positive sampling strategy for SSL frameworks, named Self-Supervised Positive Sampling (SSPS), is introduced in this work (Figure 1b). Instead of sampling a positive from the same utterance as the anchor, this bootstrapped approach determines a pseudo-positive from a different utterance by relying on the knowledge acquired and refined by the SSL model. After several epochs of standard SSL training, it is assumed that utterances of the same speaker’s identity with different recording conditions will have representations close to the anchor’s representation in the learned speaker representation space. Different sampling algorithms based on nearest neighbors (SSPS-NN) and clustering (SSPS-Clustering) are presented. This method allows learning more robust representations by matching different recordings to the same speaker identity. SSPS significantly improves the results on SV for all major SSL frameworks by decreasing the intra-class variance of speaker representations. In addition, a reduction of the information related to the source recording is observed in the learned representations when using SSPS. Furthermore, experiments show that this method is less dependent on data-augmentation as pseudo-positives are sampled from different recordings with varying channel information. This manuscript is an extended version of a previously published conference article [lepage2025SSPS], broadening the scope to a wider range of SSL frameworks, introducing two distinct sampling algorithms, and providing deeper insights into the effects of SSPS on the learned speaker representations.
The general training framework and various SSL methods are outlined in Section II. SSPS and its different sampling algorithms are then introduced in Section III. The experimental setup is detailed in Section IV, followed by a presentation of the preliminary experiments and final results in Section V. Finally, the article concludes in Section VI.
@@
Figure 2: SSL training framework for SV. The joint-embedding architecture generates a pair of representations (evaluation/downstream task) and embeddings (training/pretext task) from a given unlabeled audio sample.
@@ II Self-Supervised Learning Frameworks for Speaker Verification
The self-supervised training framework is based on a symmetrical or asymmetrical joint-embedding architecture, illustrated in Figure 2, to produce a pair of embeddings from a given unlabeled audio waveform.
Let $\mathcal{I}\equiv\{1,\ldots,N\}$ be the indices of the train set composed of $N$ samples. At each training iteration, $B$ utterances are sampled such that $\mathcal{B}\subseteq\mathcal{I}$ are the mini-batch indices. For each utterance $u_{i}\in\{u_{i}\}_{i\in\mathcal{B}}$, two frames, $\bm{x}_{i}$ and $\bm{x}_{i}^{\prime}$, are randomly extracted. They represent the anchor and the positive, respectively. Next, random data-augmentation is applied to each frame, and their mel-scaled spectrogram features are used as input features.
In the general case, the joint-embedding architecture consists of two branches, one with an encoder $f_{\theta}$ and projector $g_{\phi}$, and the other with an encoder $f_{\theta^{\prime}}$ and projector $g_{\phi^{\prime}}$, where the corresponding modules share the same architecture across branches. First, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$ transform $\bm{x}_{i}$ and $\bm{x}_{i}^{\prime}$ to their representations $\bm{y}_{i}$ and $\bm{y}_{i}^{\prime}$ of dimension $D_{\text{repr}}$. Then, projectors $g_{\phi}$ and $g_{\phi^{\prime}}$ map $\bm{y}_{i}$ and $\bm{y}_{i}^{\prime}$ to their corresponding embeddings $\bm{z}_{i}$ and $\bm{z}_{i}^{\prime}$ of dimension $D_{\text{emb}}$.
Representations are used to extract speaker representations and perform SV, while embeddings are employed to compute the loss $\mathcal{L}$ and optimize the model. For the downstream task of SV, SSL methods aim to minimize the distance between same-speaker embeddings while avoiding collapse (i.e., trivial solutions leading to non-informative representations). In this context, using distinct frames and applying data augmentation is fundamental to prevent collapse and to produce robust representations that primarily capture speaker identity.
@@
As training processes mini-batches, the following notations are introduced to represent input frames, representations, and embeddings of a mini-batch: $\bm{X}=\{\bm{x}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$, $\bm{X}^{\prime}=\{\bm{x}^{\prime}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$, $\bm{Y}=\{\bm{y}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$, $\bm{Y}^{\prime}=\{\bm{y}_{i}^{\prime}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$, $\bm{Z}=\{\bm{z}_{i}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$, and $\bm{Z}^{\prime}=\{\bm{z}_{i}^{\prime}\}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}$. $\bm{z}_{i}$ denotes the embedding of the $i$-th sample from $\bm{Z}$ and $\bm{z}^{d}$ denotes the $d$-th dimension of all embeddings in $\bm{Z}$.
In the following, five major SSL frameworks from different paradigms are presented. SimCLR and MoCo, based on contrastive learning, are described in Sections II-A and II-B, respectively. Then, a clustering-based approach, SwAV, is introduced in Section II-C. Next, VICReg, following the information maximization paradigm, is presented in Section II-D. Finally, DINO, based on self-distillation, is described in Section II-E. With this set of methods, the majority of SSL-based frameworks are covered.
The similarity between two embeddings $\bm{u}$ and $\bm{v}$ is defined as $\ell(\bm{u},\bm{v})=\exp\left(\cos\left(\bm{u},\bm{v}\right)/\tau\right)$ where $\tau$ is a temperature scaling hyperparameter. $\cos(\bm{u},\bm{v})$ corresponds to the cosine similarity and is obtained by computing the dot product between the two $l_{2}$ normalized embeddings $\bm{u}$ and $\bm{v}$.
@@ II-A SimCLR
Contrastive learning aims at maximizing the similarity within positive pairs while maximizing the distance between negative pairs. In SSL, supervision is provided by assuming that two randomly sampled utterances belong to different speakers. Positive pairs are constructed with embeddings derived from the same utterances, while negative pairs are sampled from the current mini-batch.
SimCLR [chen2020SimCLR, chen2020SimCLRV2] implements the concept of contrastive learning by optimizing $\mathcal{L}_{\text{SimCLR}}$, which is equivalent to the NT-Xent loss derived from InfoNCE [vandenoord2019CPC, rusak2025InfoNCE], defined as:
$\mathcal{L}_{\text{SimCLR}}=-\frac{1}{B}\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}\log\frac{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)}{\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}j\in\mathcal{B}}}\ell\left(\bm{z}_{i},\bm{z}_{j}^{\prime}\right)}.$ (1)
@@ II-B MoCo
MoCo [he2020MoCo, chen2020MoCoV2] follows the contrastive learning paradigm and employs an asymmetrical joint-embedding architecture, implying that the key network is updated using a momentum-based EMA of the query network weights. Additionally, a large memory queue $\bm{Q}^{\text{MoCo}}$ of recent embeddings is introduced to increase the number of negative pairs instead of sampling negatives from the current mini-batch like SimCLR. The queue is dynamically updated at each training step with embeddings of the key branch from the current mini-batch.
@@
$\mathcal{L}_{\text{MoCo}}=-\frac{1}{B}\\ \sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}i\in\mathcal{B}}}\log\frac{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)}{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)+\sum\limits_{j\in\mathcal{J}}\ell\left(\bm{z}_{i},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bm{q}^{\text{MoCo}}_{j}}\right)},$ (2)
where $\mathcal{J}\equiv\{1,\ldots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}|\bm{Q}^{\text{MoCo}}|}\}$ and $\bm{q}^{\text{MoCo}}_{i}$ is the $i$-th element of the queue $\bm{Q}^{\text{MoCo}}$.

### v3

@@ Abstract
Recent developments in Self-Supervised Learning (SSL) have demonstrated significant potential for Speaker Verification (SV), but closing the performance gap with supervised systems remains an ongoing challenge. SSL frameworks rely on anchor-positive pairs, constructed from segments of the same audio utterance. Hence, positives have channel characteristics similar to those of their corresponding anchors, even with extensive data-augmentation. Therefore, this positive sampling strategy is a fundamental limitation as it encodes too much information regarding the recording source in the learned representations. This article introduces Self-Supervised Positive Sampling (SSPS), a bootstrapped technique for sampling appropriate and diverse positives in SSL frameworks for SV. SSPS samples positives close to their anchor in the representation space, assuming that these pseudo-positives belong to the same speaker identity but correspond to different recording conditions. This method consistently demonstrates improvements in SV performance on VoxCeleb benchmarks when applied to major SSL frameworks, including SimCLR, SwAV, VICReg, and DINO. Using SSPS, SimCLR and DINO achieve 2.57% and 2.53% EER on VoxCeleb1-O, respectively. SimCLR yields a 58% relative reduction in EER, getting comparable performance to DINO with a simpler training framework. Furthermore, SSPS lowers intra-class variance and reduces channel information in speaker representations while exhibiting greater robustness without data-augmentation.
@@ I Introduction
Speaker Recognition (SR) corresponds to the process of identifying the speaker’s identity in an audio speech utterance. The main task in the SR field is Speaker Verification (SV), which aims to determine whether two speech utterances are spoken by the same speaker. To achieve this task, SR systems aim to define representations that maximize inter-speaker distances and minimize intra-speaker variances while being robust to extrinsic variabilities (e.g., noise, environment, recording device, and channel conditions).
@@
With the emergence of deep learning, traditional machine learning methods, such as i-vectors [1], have been surpassed by DNN models, including x-vectors [2], architectures based on ResNet [3], and more recently, the ECAPA-TDNN [4] model. These methods are trained to match input speech utterances to their respective speaker identities in a supervised manner on a large corpus of annotated speech segments [5]. Indeed, the performance of deep learning methods scales with the amount of training data. However, this training paradigm represents a notable constraint due to the scarcity and expense of getting annotated training samples.
@@
Self-Supervised Learning (SSL) methods have emerged as a solution to this bottleneck by learning meaningful representations from the input data. SSL has been shown to enhance the scalability potential of models by leveraging the abundance of unlabeled data available. Various SSL frameworks have been proposed in computer vision following the joint-embedding architecture, which considers an anchor and a positive sample derived from different augmentations of the same input, thereby representing the same high-level information. Contrastive learning, adopted by SimCLR [6] and MoCo [7], aims to maximize the similarity of positive pairs while minimizing the similarity of negative pairs sampled from the current batch or a memory queue. The effect of class collision, occurring when randomly sampled negatives belong to the same class as the anchor-positive pair, is considered negligible on the training convergence. Clustering-based methods, such as DeepCluster [8] and SwAV [9], learn by grouping similar representations into clusters and using the resulting assignments as a supervisory signal. Information maximization techniques like Barlow Twins [10] and VICReg [11] apply constraints on the learned embedding space to enforce invariance between the anchor and the positive representations while explicitly avoiding collapse (i.e., non-informative representations). Finally, methods based on self-distillation, such as BYOL [12] and DINO [13], leverage knowledge distillation, where a student model is trained to match the output of a teacher model.
@@
These approaches have also been successfully extended to the downstream task of SV. Most methods are based on contrastive learning [14, 15, 16, 17] and self-distillation, as the DINO framework currently achieves state-of-the-art performance on SV [18, 19, 20, 21, 22]. In this context, SSL relies on the assumption that the anchor and the positive belong to the same speaker, as both segments are extracted from the same speech utterance. Moreover, extensive data-augmentation is applied to both inputs to learn representations robust against extrinsic variabilities (e.g., environmental noise, reverberation, acoustic conditions, recording devices).
@@
Nonetheless, data-augmentation is insufficient to overcome the effect of SSL same-utterance positive sampling, which encodes various channel characteristics in speaker representations, yielding high variance within same-speaker representations. Several techniques have been explored to address this issue in SV: AP+AAT [14] introduces an adversarial loss to penalize the model from learning channel information; i-mix [23] is a data-driven augmentation strategy that interpolates training utterances to make the model focus on the main distinguishing factor between them; DPP [24] finds diverse positive samples using another modality by cross-referencing speech and face data; CA-DINO [20] clusters the embeddings to sample a positive from the same speaker class as the anchor. Finally, alternative positive sampling strategies have been proposed in computer vision, such as NNCLR [25] and GPS-SSL [26], which select positives in the latent space through nearest neighbor search.
@@
A new positive sampling strategy for SSL frameworks, named Self-Supervised Positive Sampling (SSPS), is introduced in this article. Instead of sampling a positive from the same utterance as the anchor (Figure LABEL:sub@fig:ssps_overview_ssl), this bootstrapped approach determines a pseudo-positive from a different utterance by relying on the knowledge acquired and refined by the SSL model (Figure LABEL:sub@fig:ssps_overview_ssps). After several epochs of standard SSL training, it is assumed that utterances of the same speaker identity with different recording conditions will have representations close to the anchor’s representation in the learned speaker representation space. Different sampling algorithms based on nearest neighbors (SSPS-NN) and clustering (SSPS-Clustering) are presented. This method enables the learning of more robust representations by matching different recordings to the same speaker identity. SSPS significantly improves the performance on SV for all major SSL frameworks by reducing the intra-class variance of speaker representations. Additionally, a decrease in the information related to the source recording is observed in the learned representations when using SSPS. Furthermore, experiments show that this method is less dependent on data-augmentation as pseudo-positives are sampled from different recordings with varying channel information. This manuscript is an extended version of a previously published conference article [27], broadening the scope to a wider range of SSL frameworks, introducing two distinct sampling algorithms, and providing deeper insights into the effects of SSPS on the learned speaker representations.
The general SSL training scheme and major frameworks are described in Section II. Then, SSPS and its instantiations, SSPS-NN and SSPS-Clustering, are introduced in Section III. The experimental setup is detailed in Section IV. Subsequently, Section V highlights the limitations of SSL positive sampling and presents preliminary results obtained with SSPS. It then evaluates SSPS on SV, compares it with state-of-the-art methods, and analyzes the resulting learned representations. Finally, the article concludes in Section VI.
@@
Figure 2: SSL training framework for SV. The joint-embedding architecture generates a pair of representations (for the downstream task) and a pair of embeddings (for the pretext task) from a given unlabeled audio sample.
@@ II Self-Supervised Learning Frameworks for Speaker Verification
The self-supervised training framework is based on the symmetrical or asymmetrical joint-embedding architecture to produce a pair of embeddings from a given unlabeled audio waveform, as illustrated in Figure 2.
Let $\mathcal{I}\equiv\{1,\ldots,N\}$ be the indices of the training set composed of $N$ samples. At each training iteration, $B$ utterances are sampled such that $\mathcal{B}\subseteq\mathcal{I}$ are the mini-batch indices. For each utterance $u_{i}\in\{u_{i}\}_{i\in\mathcal{B}}$, two segments, $\bm{x}_{i}$ and $\bm{x}_{i}^{\prime}$, are randomly extracted. They represent the anchor and the positive, respectively. Next, random data-augmentation is applied to each segment, and their mel-scaled spectrogram features are used as input features.
The joint-embedding architecture consists of two branches, one with an encoder $f_{\theta}$ and projector $g_{\phi}$, and the other with an encoder $f_{\theta^{\prime}}$ and projector $g_{\phi^{\prime}}$, where the corresponding modules share the same architecture across branches. First, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$ transform $\bm{x}_{i}$ and $\bm{x}_{i}^{\prime}$ to their representations $\bm{y}_{i}$ and $\bm{y}_{i}^{\prime}$ of dimension $D_{\text{repr}}$. Then, projectors $g_{\phi}$ and $g_{\phi^{\prime}}$ map $\bm{y}_{i}$ and $\bm{y}_{i}^{\prime}$ to their corresponding embeddings $\bm{z}_{i}$ and $\bm{z}_{i}^{\prime}$ of dimension $D_{\text{emb}}$.
Representations serve as speaker features to perform SV, while embeddings are used to compute the loss $\mathcal{L}$ and optimize the model. For the downstream task of SV, SSL methods aim to minimize the distance between same-speaker embeddings while avoiding collapse (i.e., trivial solutions leading to non-informative representations). In this context, using distinct segments and applying data-augmentation is fundamental to prevent collapse and to produce robust representations that primarily capture speaker identity.
@@
As training processes mini-batches, the following notations are introduced to represent input segments, representations, and embeddings: $\bm{X}=\{\bm{x}_{i}\}_{i\in\mathcal{B}}$, $\bm{X}^{\prime}=\{\bm{x}^{\prime}_{i}\}_{i\in\mathcal{B}}$, $\bm{Y}=\{\bm{y}_{i}\}_{i\in\mathcal{B}}$, $\bm{Y}^{\prime}=\{\bm{y}_{i}^{\prime}\}_{i\in\mathcal{B}}$, $\bm{Z}=\{\bm{z}_{i}\}_{i\in\mathcal{B}}$, and $\bm{Z}^{\prime}=\{\bm{z}_{i}^{\prime}\}_{i\in\mathcal{B}}$. $\bm{z}_{i}$ denotes the embedding of the $i$-th sample from $\bm{Z}$ and $\bm{z}^{d}$ denotes the $d$-th dimension of all embeddings in $\bm{Z}$.
The following presents major SSL frameworks from different paradigms. SimCLR and MoCo, based on contrastive learning, are described in Sections II-A and II-B, respectively. Then, SwAV, a clustering-based approach, is introduced in Section II-C. Next, VICReg, following the information maximization paradigm, is presented in Section II-D. Finally, DINO, based on self-distillation, is described in Section II-E. This set of methods allows the majority of SSL-based frameworks to be covered.
The similarity between two embeddings $\bm{u}$ and $\bm{v}$ is defined as $\ell(\bm{u},\bm{v})=\exp\left(\operatorname{sim}\left(\bm{u},\bm{v}\right)/\tau\right)$ where $\tau$ is a temperature scaling hyperparameter. $\operatorname{sim}(\bm{u},\bm{v})$ corresponds to the cosine similarity and is obtained by computing the dot product between the two $l_{2}$ normalized embeddings $\bm{u}$ and $\bm{v}$.
@@ II-A SimCLR
SimCLR [6, 28] implements the concept of contrastive learning, which aims to maximize the similarity between the anchor and the positive, while minimizing the similarity between the anchor and the negative samples. Negatives are randomly selected from the current mini-batch, assuming that these utterances belong to a different speaker identity.
The objective function $\mathcal{L}_{\text{SimCLR}}$, which is equivalent to the NT-Xent loss derived from InfoNCE [29, 30], is defined as:
$\mathcal{L}_{\text{SimCLR}}=-\frac{1}{B}\sum_{i\in\mathcal{B}}\log\frac{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)}{\sum_{j\in\mathcal{B}}\ell\left(\bm{z}_{i},\bm{z}_{j}^{\prime}\right)}.$ (1)
@@ II-B MoCo
MoCo [7, 31] follows the contrastive learning paradigm and employs the asymmetrical joint-embedding architecture, implying that the key weights are updated with an EMA of the query weights. A queue $\bm{Q}^{\text{MoCo}}$ of recent key embeddings, dynamically updated at each training step, is introduced to increase the number of negatives.
@@
$\mathcal{L}_{\text{MoCo}}=-\frac{1}{B}\\ \sum_{i\in\mathcal{B}}\log\frac{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)}{\ell\left(\bm{z}_{i},\bm{z}_{i}^{\prime}\right)+\sum\limits_{j\in\mathcal{J}}\ell\left(\bm{z}_{i},\bm{q}^{\text{MoCo}}_{j}\right)},$ (2)
where $\mathcal{J}\equiv\{1,\ldots,|\bm{Q}^{\text{MoCo}}|\}$ and $\bm{q}^{\text{MoCo}}_{i}$ is the $i$-th element of the queue $\bm{Q}^{\text{MoCo}}$.